### v1

# Hadronic tensor molecule $B_{c}^{\ast+}B_{c}^{\ast-}$

Abstract
Mass and full decay width of the hadronic tensor molecule $\mathcal{M}_{\mathrm{T}}=B_{c}^{\ast+}B_{c}^{\ast-}$ are examined in the framework of QCD sum rule method. To find the mass $m$ and current coupling $\Lambda$ of this state, we use the two-point sum rules. The result $(12.87\pm 0.08)~\mathrm{GeV}$ for $m$ indicates that $\mathcal{M}_{\mathrm{T}}$ can easily decay to pairs of $J/\psi\Upsilon$, $\eta_{b}\eta_{c}$, $B_{c}^{+}B_{c}^{-}$, and $B_{c}^{\ast+}B_{c}^{\ast-}$ mesons. Kinematically permitted ways for $\mathcal{M}_{\mathrm{T}}$ to transform to conventional mesons include also processes $\mathcal{M}_{\mathrm{T}}\to D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$, and $D_{s}^{(\ast)+}D_{s}^{(\ast)-}$ triggered by annihilation of $b\overline{b}$ quarks in the molecule. The decays to meson pairs $B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{(\ast)0}\overline{B}_{s}^{(\ast)0}$ generated by $c\overline{c}\to$ light quarks are also among possible channels of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$. Technical tools of the three-point sum rule approach are applied to compute partial widths of these decays. Our results for the mass and width $\Gamma_{\mathcal{M}_{\mathrm{T}}}=(154\pm 19)~\mathrm{MeV}$ of the tensor molecule $\mathcal{M}_{\mathrm{T}}$ are important for experimental studies of fully heavy exotic structures.
I Introduction
Experimental discoveries of new four $X$ structures reported by the LHCb, ATLAS and CMS Collaborations had important consequences for physics of exotic four-quark mesons composed of exclusively heavy quarks LHCb:2020bwg ; Bouhova-Thacker:2022vnt ; CMS:2023owd . These $X$ states have the masses 6.27.3GeV6.2-7.3\ $\mathrm${GeV} and are presumably scalar resonances with the content $cc\overline{c}\overline{c}$. Note that their interpretations as kinematical effects were suggested as well.
Exotic mesons composed of four heavy quarks were studied during long time as interesting structures. The reason is that neither the quark-parton model nor quantum chromodynamics (QCD) forbid existence of states beyond the $q\overline{q}$ and $qq^{\prime}q^{\prime\prime}$ spectroscopy. Experimental observations of recent years supported interest to such unconventional hadrons.
The tetraquarks $bc\overline{b}\overline{c}$ with various quantum numbers are such particles. It is possible that they will be soon observed in experiments Carvalho:2015nqf ; Abreu:2023wwg . These tetraquarks may have different internal organization. First, the colored diquark $bc$ and antidiquark $\overline{b}\overline{c}$ may form a colorless diquark-antidiquark state with various quantum numbers. Second, $\overline{c}b$ and $b\overline{c}$ establish conventional $B_{c}$ mesons, and a tetraquark emerges as the hadronic $B_{c}B_{c}$ molecule state.
The diquark-antidiquark structures $bc\overline{b}\overline{c}$ with numerous quantum numbers were intensively investigated in the literature Faustov:2022mvs ; Wu:2016vtq ; Liu:2019zuc ; Chen:2019vrj ; Bedolla:2019zwg ; Cordillo:2020sgc ; Weng:2020jao ; Yang:2021zrc ; Hoffer:2024alv . In these articles the authors concentrated mainly on their masses and computed this parameter by employing different methods. It should be noted that the decays of the tetraquarks $bc\overline{b}\overline{c}$, as usual, remained beyond the scope of these studies. In our publications Agaev:2024wvp ; Agaev:2024mng ; Agaev:2024qbh we considered the diquark-antidiquark states $bc\overline{b}\overline{c}$ with spin-parities $J^{\mathrm{P}}=0^{+}$, $1^{+}$, $2^{+}$ and calculated their masses and widths.
The molecular structures $B_{c}B_{c}$ were objects of explorations as well Liu:2023gla ; Liu:2024pio ; Agaev:2025wdj ; Agaev:2025fwm . In Refs. Liu:2023gla ; Liu:2024pio they were studied in the framework of the coupled-channel unitary approach. The masses and widths of the scalar $B_{c}^{+}B_{c}^{-}$ and axial-vector $(B_{c}^{\ast+}B_{c}^{-}+B_{c}^{+}B_{c}^{\ast-})/2$ molecules were calculated in our works as well Agaev:2025wdj ; Agaev:2025fwm . To find parameters of these molecular states, we applied QCD sum rule (SR) method Shifman:1978bx ; Shifman:1978by . Our predictions for the masses $(12.725\pm 0.085)~\mathrm{GeV}$ and $(12.77\pm 0.06)~\mathrm{GeV}$ of these compounds demonstrated that they are unstable against strong decays into ordinary mesons. Their full decay widths $(155\pm 23)~\mathrm{MeV}$ and $(93\pm 14)~\mathrm{MeV}$, respectively, characterize them as relatively broad exotic mesons.
In this article, we continue our studies of $B_{c}$ mesons’ molecules and explore the hadronic tensor state $\mathcal{M}_{\mathrm{T}}=B_{c}^{\ast+}B_{c}^{\ast-}$. We are going to calculate its mass $m$ and current coupling $\Lambda$ (pole residue) using the two-point sum rule method. The three-point sum rule approach is applied to treat numerous decay channels of this state. Technical tools of this approach are required to evaluate the form factors that describe strong interaction of particles at $\mathcal{M}_{\mathrm{T}}M_{1}M_{2}$ vertices, where $M_{1}$ and $M_{2}$ are final-state mesons with appropriate charges and spin-parities. In its turns, these form factors at the mass shell of one of the mesons allow us to estimate the strong couplings at $\mathcal{M}_{\mathrm{T}}M_{1}M_{2}$.
Note that there are different mechanisms that trigger decay of the molecule $\mathcal{M}_{\mathrm{T}}$. In fact, $\mathcal{M}_{\mathrm{T}}$ can decay to heavy mesons $J/\psi\Upsilon$ and $\eta_{b}\eta_{c}$, $B_{c}^{\ast+}B_{c}^{\ast-}$ and $B_{c}^{+}B_{c}^{-}$, in which all constituent quarks of the hadronic molecule participate in forming of the final-state particles. These processes are dominant decay modes of $\mathcal{M}_{\mathrm{T}}$. Another type of decays are due to subprocesses $b\overline{b}$ or $c\overline{c}\rightarrow$ $q\overline{q}$, $s\overline{s}$ inside of $\mathcal{M}_{\mathrm{T}}$ with subsequent creation of $D_{(s)}D_{(s)}$ and $B_{(s)}B_{(s)}$ pairs provided the mass $m$ and quantum numbers of $D_{(s)}$ and $B_{(s)}$ mesons make such processes kinematically possible Becchi:2020mjz ; Becchi:2020uvq ; Agaev:2023ara .
The current paper is divided into seven sections: In Sec. II, we compute the spectroscopic parameters of the hadronic tensor molecule $\mathcal{M}_{\mathrm{T}}$. The decays $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$ and $\eta_{b}\eta_{c}$ are considered in Sec. III, in which we evaluate their partial widths. The decays into $B_{c}^{\ast+}B_{c}^{\ast-}$ and $B_{c}^{+}B_{c}^{-}$ mesons are explored in Sec. IV. The widths of the processes $\mathcal{M}_{\mathrm{T}}\rightarrow D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$ and $D_{s}^{(\ast)+}D_{s}^{(\ast)-}$ are examined in Sec. V. Section VI is devoted to analysis of $\mathcal{M}_{\mathrm{T}}$ molecule’s decays to $B$ meson pairs, i.e. to decays $B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{0}\overline{B}_{s}^{0}.$ Here, we also evaluate the full width of the molecule $\mathcal{M}_{\mathrm{T}}$. Discussion of obtained results and related conclusions are made in Sec. VII.
II Spectroscopic parameters $m$ and $\Lambda$ of the molecule $\mathcal{M}_{\mathrm{T}}$
The mass $m$ and current coupling (pole residue) $\Lambda$ of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$ are important quantities that determine properties of this state. One of the powerful nonperturbative approaches to determine them is the two-point sum rule method. In this context we have to derive the SRs for these parameters. For these purposes, we analyze the correlation function
$\Pi_{\mu\nu\alpha\beta}(p)=i\int d^{4}xe^{ipx}\langle 0|\mathcal{T}\{J_{\mu\nu}(x)J_{\alpha\beta}^{{\dagger}}(0)\}|0\rangle,$ (1)
where $J_{\mu\nu}(x)$ is the interpolating current for the tensor molecule $\mathcal{M}_{\mathrm{T}}$. The time-ordered product of two currents in Eq. (1) is denoted by $\mathcal{T}$.
The current $J_{\mu\nu}(x)$ for the molecule $\mathcal{M}_{\mathrm{T}}=B_{c}^{\ast+}B_{c}^{\ast-}$ has the form
$J_{\mu\nu}(x)=[\overline{b}_{a}(x)\gamma_{\mu}c_{a}(x)][\overline{c}_{b}(x)\gamma_{\nu}b_{b}(x)],$ (2)
where $a$ and $b$ are the color indices. The current $J_{\mu\nu}$ corresponds to the state with quantum numbers $J^{\mathrm{PC}}=2^{++}$.
The SRs for $m$ and $\Lambda$ can be obtained by equating $\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p)$ and $\Pi_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p)$. The first function $\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p)$ is the correlator expressed using the physical parameters of the molecule $\mathcal{M}_{\mathrm{T}}$. To find it, one inserts into Eq. (1) a full set of states with the quark content and quantum numbers of $\mathcal{M}_{\mathrm{T}}$, and carries out integration over $x$. Finally, $\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p)$ becomes equal to
$\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p)$ $=$ $\frac{\langle 0|J_{\mu\nu}|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle\langle\mathcal{M}_{\mathrm{T}}(p,\epsilon)|J_{\alpha\beta}^{{\dagger}}|0\rangle}{m^{2}-p^{2}}$ (3)
$+\cdots.$
Here, the term shown in Eq. (3) is due to the ground-level molecule $\mathcal{M}_{\mathrm{T}}$, while the ellipses denote contributions of higher resonances and continuum states. To proceed with computations, we use the matrix element
$\langle 0|J_{\mu\nu}|\mathcal{M}_{\mathrm{T}}(p,\epsilon(p)\rangle=\Lambda\epsilon_{\mu\nu}^{(\lambda)}(p),$ (4)
where $\epsilon=\epsilon_{\mu\nu}^{(\lambda)}(p)$ is the polarization tensor of the molecule $\mathcal{M}_{\mathrm{T}}$. We replace in $\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p)$ matrix elements in the correlator Eq. (3) by Eq. (4) and perform required operations. Our computations yield
$\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p)$ $=$ $\frac{\Lambda^{2}}{m^{2}-p^{2}}\left\{\frac{1}{2}\left(g_{\mu\alpha}g_{\nu\beta}+g_{\mu\beta}g_{\nu\alpha}\right)\right$. (5)
+ $\left.+\text{ other terms}\right\}+\cdots,$
where dots stand for higher resonances and continuum states. Note that there are various Lorentz structures in Eq. (5). The component $\sim(g_{\mu\alpha}g_{\nu\beta}+g_{\mu\beta}g_{\nu\alpha})$ is formed due to only the spin-$2$ particle. Therefore, we choose for our studies this component of the correlation function and denote by $\Pi^{\mathrm{Phys}}(p^{2})$ the related invariant amplitude.
To compute $\Pi_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p)$ one inserts $J_{\mu\nu}(x)$ into Eq. (1) and contract quark fields. In the case under discussion, we get
$\Pi_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p)=i\int d^{4}xe^{ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{aa^{\prime}}(x)\gamma_{\beta}S_{b}^{a^{\prime}a}(-x)\right]$
$\times\mathrm{Tr}\left[\gamma_{\nu}S_{b}^{bb^{\prime}}(x)\gamma_{\alpha}S_{c}^{b^{\prime}b}(-x)\right],$ (6)
where $S_{b(c)}(x)$ are heavy-quark propagators Agaev:2020zad . The correlator $\Pi_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p)$ obtained by this way has to be calculated in the operator product expansion ($\mathrm{OPE}$) with certain accuracy.
Having extracted the structure $(g_{\mu\alpha}g_{\nu\beta}+g_{\mu\beta}g_{\nu\alpha})$ from $\Pi_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p)$ and labeled the relevant invariant amplitude by $\Pi^{\mathrm{OPE}}(p^{2})$, one finds SRs for $m$ and $\Lambda$, which reads
$m^{2}=\frac{\Pi^{\prime}(M^{2},s_{0})}{\Pi(M^{2},s_{0})},$ (7)
and
$\Lambda^{2}=e^{m^{2}/M^{2}}\Pi(M^{2},s_{0}).$ (8)
The function $\Pi(M^{2},s_{0})$ is the amplitude $\Pi^{\mathrm{OPE}}(p^{2})$ obtained after the Borel transformation and continuum subtraction, whereas $\Pi^{\prime}(M^{2},s_{0})$ is its derivative over $d/d(-1/M^{2})$. The Borel transformation is applied to suppress contribution of higher resonances and continuum states, whereas these terms are subtracted from the $\Pi(M^{2},\infty)$ using ideas of the quark-hadron duality and continuum threshold parameter $s_{0}$. As a result, $\Pi(M^{2},s_{0})$ depends on $s_{0}$ and on the Borel parameter $M^{2}$. It has the following form.
$\Pi(M^{2},s_{0})=\int_{4(m_{b}+m_{c})^{2}}^{s_{0}}ds\rho^{\mathrm{OPE}}(s)e^{-s/M^{2}}+\Pi(M^{2}).$ (9)
The spectral density $\rho^{\mathrm{OPE}}(s)$ is calculated as the imaginary piece of the amplitude $\Pi^{\mathrm{OPE}}(p^{2})$. The nonperturbative contribution $\Pi(M^{2})$ is found straightly from $\Pi^{\mathrm{OPE}}(p)$ and encompasses terms do not included into $\rho^{\mathrm{OPE}}(s)$.
To conduct numerical calculations one needs to choose input parameters in Eqs. (7) and (8). The masses $m_{c}$ and $m_{b}$ of the quarks as well as the condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ are universal quantities. In the current article, we employ
$m_{c}=(1.2730\pm 0.0046)~\mathrm{GeV},$
$m_{b}=(4.183\pm 0.007)~\mathrm{GeV},$
$\langle\alpha_{s}G^{2}/\pi\rangle=(0.012\pm 0.004)~\mathrm{GeV}^{4}.$ (10)
The quark masses $m_{c}$ and $m_{b}$ are calculated in the $\overline{\mathrm{MS}}$ scheme PDG:2024 . The condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ was estimated in Refs. Shifman:1978bx ; Shifman:1978by from studies of different processes.
The Borel $M^{2}$ and continuum subtraction $s_{0}$ parameters depend on the process under analysis and should satisfy the usual restrictions of the sum rule calculations. Prevalence of the pole contribution ($\mathrm{PC}$) in the mass $m$ and current coupling $\Lambda$, their independence on $M^{2}$ and $s_{0}$, as well as convergence of $\mathrm{OPE}$ are decisive for the credible SR analysis. To obey these constraints, we impose on $M^{2}$ and $s_{0}$ the following requirements. First, the pole contribution
$\mathrm{PC}=\frac{\Pi(M^{2},s_{0})}{\Pi(M^{2},\infty)},$ (11)
has to satisfy $\mathrm{PC}\geq 0.5$ which leads to its dominance in $m$ and $\Lambda$. The correlation function contains only the dimension-$4$ term $\Pi^{\mathrm{Dim4}}(M^{2},s_{0})$, therefore we demand fulfilment of the constraint $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$: This ensures the convergence of $\mathrm{OPE}$. The maximum of the Borel parameter is fixed by Eq. (11), whereas convergence of $\mathrm{OPE}$ permits us to obtain its minimal value.
Figure 1: The $\mathrm{PC}$ as a function of $M^{2}$ at various $s_{0}$. The border $\mathrm{PC}=0.5$ is plotted by the horizontal line. The red triangle fixes the point $M^{2}=9.5~\mathrm{GeV}^{2}$ and $s_{0}=185~\mathrm{GeV}^{2}$.
Figure 2: Mass $m$ as a function of the Borel $M^{2}$ (left panel), and continuum threshold $s_{0}$ parameters (right panel).
Numerical computations are carried out for different values of $M^{2}$ and $s_{0}$. Examination of obtained results permits us to determine regions for $M^{2}$ and $s_{0}$, where all restrictions are satisfied. We find that the regions
$M^{2}\in[8.5,10.5]~\mathrm{GeV}^{2},\ s_{0}\in[184,186]~\mathrm{GeV}^{2}$ (12)
comply with these constraints. In fact, at maximal and minimal values of $M^{2}$ the pole contribution on the average in $s_{0}$ is $\mathrm{PC}\approx 0.50$ and $\mathrm{PC}\approx 0.68$, correspondingly. The nonperturbative term is negative and establishes at $M^{2}=9.5~\mathrm{GeV}^{2}$ approximately $1\%$ of the full result. The parameter $\mathrm{PC}$ as a function of $M^{2}$ is depicted in Fig. 1, where all curves overshoot the border $\mathrm{PC}=0.5$.
We compute mean values of $m$ and $\Lambda$ over the regions Eq. (12) and find
$m=(12.87\pm 0.08)~\mathrm{GeV},$
$\Lambda=(6.70\pm 0.78)\times 10^{-2}~\mathrm{GeV}^{5}.$ (13)
The results in Eq. (13) amount to SR predictions at the point $M^{2}=9.5~\mathrm{GeV}^{2}$ and $s_{0}=185~\mathrm{GeV}^{2}$, where $\mathrm{PC}\approx 0.59$. This guaranties the prevalence of $\mathrm{PC}$ in the quantities $m$ and $\Lambda$. The main sources of ambiguities in Eq. (13) are parameters $M^{2}$ and $s_{0}$. Relevant errors constitutes $\pm 0.6\%$ of the mass $m$, which proves reliability of the extracted result. Such accuracy is achieved due to the sum rule for $m$, Eq. (7), which is a ratio of correlators. Consequently, variations of these functions due to changes of $M^{2}$, and $s_{0}$ recompense in $m$ and stabilize the result. Ambiguities of the current coupling $\Lambda$ are equal to $\pm 11.6\%$ remaining nevertheless inside borders allowable for SR studies. In Fig. 2, we depict $m$ as a function of $M^{2}$ and $s_{0}$.
III Processes $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$, $\eta_{b}\eta_{c}$
Parameters of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$ fix its kinematically allowed decay modes. Decays to mesons $J/\psi\Upsilon$ and $\eta_{b}\eta_{c}$ are such channels. In fact, thresholds for these processes are $12.558~\mathrm{GeV}$ and $12.383~\mathrm{GeV}$, respectively. Here, we concentrate on the decays $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$ and $\mathcal{M}_{\mathrm{T}}\rightarrow\eta_{b}\eta_{c}$.
III.1 Decay $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$
First, we explore the channel $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$ and compute its partial width. For these purposes, we should determine the strong coupling $g_{1}$ at the vertex $\mathcal{M}_{\mathrm{T}}J/\psi\Upsilon$. This can be done by evaluating the form factor $g_{1}(q^{2})$ at the mass shell $q^{2}=m_{J/\psi}^{2}$.
We extract SR for the form factor $g_{1}(q^{2})$ from analysis of the three-point correlator
$\Pi_{\mu\nu\alpha\beta}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{\Upsilon}(y)$ (14)
$\times J_{\nu}^{J/\psi}(0)J_{\alpha\beta}^{\dagger}(x)\}|0\rangle,$
with $J_{\mu}^{\Upsilon}(x)$ and $J_{\nu}^{J/\psi}(x)$ being the interpolating currents of the vector quarkonia $\Upsilon$ and $J/\psi$, respectively. They are given by expressions
$J_{\mu}^{\Upsilon}(x)=\overline{b}_{i}(x)\gamma_{\mu}b_{i}(x),\ J_{\nu}^{J/\psi}(x)=\overline{c}_{j}(x)\gamma_{\nu}c_{j}(x),$ (15)
where $i$ and $j$ are the color indices.
The physical side of the sum rule $\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$ is found by rewriting Eq. (14) using parameters of the particles $\mathcal{M}_{\mathrm{T}}$, $J/\psi$, and $\Upsilon$. By considering only the ground-level states, we convert $\Pi_{\mu\nu\alpha\beta}(p,p^{\prime})$ into the form
$\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{\Upsilon}|\Upsilon(p^{\prime},\varepsilon_{1})\rangle}{p^{\prime 2}-m_{\Upsilon}^{2}}\frac{\langle 0|J_{\nu}^{J/\psi}|J/\psi(q,\varepsilon_{2})\rangle}{q^{2}-m_{J/\psi}^{2}}$
$\times\langle\Upsilon(p^{\prime},\varepsilon_{1})J/\psi(q,\varepsilon_{2})|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle\frac{\langle\mathcal{M}_{\mathrm{T}}(p,\varepsilon)|J_{\alpha\beta}^{\dagger}|0\rangle}{p^{2}-m^{2}}$
$+\cdots,$ (16)
where $m_{\Upsilon}=(9460.40\pm 0.10)~\mathrm{MeV}$ and $m_{J/\psi}=(3096.900\pm 0.006)~\mathrm{MeV}$ are masses of the $\Upsilon$ and $J/\psi$ mesons PDG:2024 . Above, $\varepsilon_{1}$ and $\varepsilon_{2}$ are the polarization vectors of the mesons $\Upsilon$ and $J/\psi$, respectively.
Equation (16) can be written in a more simple form. To this end, we employ the matrix elements
$\langle 0|J_{\mu}^{\Upsilon}|\Upsilon(p^{\prime},\varepsilon_{1})\rangle$ $=$ $f_{\Upsilon}m_{\Upsilon}\varepsilon_{1\mu}(p^{\prime}),$
$\langle 0|J_{\nu}^{J/\psi}|J/\psi(q,\varepsilon_{2})\rangle$ $=$ $f_{J/\psi}m_{J/\psi}\varepsilon_{2\nu}(q).$ (17)
Here, $f_{\Upsilon}=(708\pm 8)~\mathrm{MeV}$ and $f_{J/\psi}=(411\pm 7)~\mathrm{MeV}$ are the experimental values of the mesons’ decay constants Lakhina:2006vg .
We also have to specify the matrix element $\langle\Upsilon(p^{\prime},\varepsilon_{1})J/\psi(q,\varepsilon_{2})|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle$ which can be done by expressing it in terms of parameters of the particles $\mathcal{M}_{\mathrm{T}}$, $\Upsilon$ and $J/\psi$ and corresponding form factors. Further analysis demonstrates that the tensor-vector-vector vertex has the form Agaev:2024pil
$\langle\Upsilon(p^{\prime},\varepsilon_{1})J/\psi(q,\varepsilon_{2})|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle=g_{1}(q^{2})\epsilon_{\tau\rho}^{(\lambda)}\left[(\varepsilon_{1}^{\ast}\cdot q)\right$.
$\left.\times\varepsilon_{2}^{\tau\ast}p^{\prime\rho}+(\varepsilon_{2}^{\ast}\cdot p^{\prime})\varepsilon_{1}^{\ast\tau}q^{\rho}-(p^{\prime}\cdot q)\varepsilon_{1}^{\tau\ast}\varepsilon_{2}^{\rho\ast}-(\varepsilon_{1}^{\ast}\cdot\varepsilon_{2}^{\ast})p^{\prime\tau}q^{\rho}\right].$
Then, for $\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$ we get
$\Pi_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})=g_{1}(q^{2})\frac{\Lambda f_{\Upsilon}m_{\Upsilon}f_{J/\psi}m_{J/\psi}}{\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{\Upsilon}^{2})(q^{2}-m_{J/\psi}^{2})}$
$\times\left[p_{\beta}^{\prime}p_{\alpha}^{\prime}g_{\mu\nu}+\frac{1}{2}p_{\mu}p_{\alpha}^{\prime}g_{\beta\nu}+\frac{1}{2m^{2}}p_{\beta}p_{\nu}p_{\mu}^{\prime}p_{\alpha}^{\prime}\right$.
+ $\left.+\text{ other structures}\right]+\cdots.$ (19)
The correlation function $\Pi_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p,p^{\prime})$ is given by the formula
$\Pi_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p,p^{\prime})=\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ib}(y-x)\right$.
$\left.\times\gamma_{\alpha}S_{c}^{bj}(x)\gamma_{\nu}S_{c}^{ja}(-x)\gamma_{\beta}S_{b}^{ai}(x-y)\right].$ (20)
We use the amplitudes $\Pi_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ and $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ that correspond to terms $\sim p_{\mu}p_{\alpha}^{\prime}g_{\beta\nu}$ in these correlators, and determine SR for $g_{1}(q^{2})$. After standard manipulations, we get
$g_{1}(q^{2})=\frac{2(q^{2}-m_{J/\psi}^{2})}{\Lambda f_{\Upsilon}m_{\Upsilon}f_{J/\psi}m_{J/\psi}}e^{m^{2}/M_{1}^{2}}e^{m_{\Upsilon}^{2}/M_{2}^{2}}\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (21)
In Eq. (21), $\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ is the function $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ after the Borel transformations and continuum subtractions. As is clear it contains $\mathbf{M}^{2}=(M_{1}^{2},M_{2}^{2})$ and $\mathbf{s}_{0}=(s_{0},s_{0}^{\prime})$ where the pairs $(M_{1}^{2},s_{0})$ and $(M_{2}^{2},s_{0}^{\prime})$ correspond to the molecule and $\Upsilon$ channels, respectively. It is given by the formula
$\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})=\int_{4\mathcal{M}^{2}}^{s_{0}}ds\int_{4m_{b}^{2}}^{s_{0}^{\prime}}ds^{\prime}\rho_{1}(s,s^{\prime},q^{2})$
$\times e^{-s/M_{1}^{2}-s^{\prime}/M_{2}^{2}}.$ (22)
Constraints used to fix the quantities $\mathbf{M}^{2}$ and $\mathbf{s}_{0}$ are standard for all sum rule investigations and have been outlined in Sec. II. Our calculations show that the regions in Eq. (12) for the parameters $(M_{1}^{2},s_{0})$ and
$M_{2}^{2}\in[10,12]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[98,100]~\mathrm{GeV}^{2}.$ (23)
for $(M_{2}^{2},s_{0}^{\prime})$ satisfy these requirements. Let us note that $s_{0}^{\prime}$ is bounded by the mass $m_{\Upsilon(2S)}=(10023.4\pm 0.5)~\mathrm{MeV}\$of the excited state $\Upsilon(2S)$, i.e., $s_{0}^{\prime}<m_{\Upsilon(2S)}^{2}$.
The sum rule gives reliable predictions for the form factor $g_{1}(q^{2})$ in the Euclidean region $q^{2}<0$. But $g_{1}(q^{2})$ determines the coupling $g_{1}$ at the mass shell $q^{2}=m_{J/\psi}^{2}$. Therefore, it is convenient to employ the function $g_{1}(Q^{2})$ with $Q^{2}=-q^{2}$ and use it in our analysis. The SR results for $g_{1}(Q^{2})$ are plotted in Fig. 3, where $Q^{2}$ varies in the region $Q^{2}=2-30~\mathrm{GeV}^{2}$.
Above it has been noted that $g_{1}$ should be extracted at $q^{2}=m_{J/\psi}^{2}$, i.e., at $Q^{2}=-m_{J/\psi}^{2}$ where the SR method can not be applied. Therefore, we use the function $\mathcal{G}_{1}(Q^{2})$ which at $Q^{2}>0$ is equal to the SR data $g_{1}(Q^{2})$, but can be extrapolated to the region $Q^{2}<0$. To this end, we introduce the function
$\mathcal{G}_{i}(Q^{2})=\mathcal{G}_{i}^{0}\mathrm{\exp}\left[c_{i}^{1}\frac{Q^{2}}{m^{2}}+c_{i}^{2}\left(\frac{Q^{2}}{m^{2}}\right)^{2}\right],$ (24)
where $\mathcal{G}_{i}^{0}$, $c_{i}^{1}$, and $c_{i}^{2}$ are fitted constants. From confronting of the SR data and Eq. (24), it is not difficult to find
$\mathcal{G}_{1}^{0}=0.32~\mathrm{GeV}^{-1},c_{1}^{1}=2.13,\text{ }c_{1}^{2}=1.34.$ (25)
The function $\mathcal{G}_{1}(Q^{2})$ is plotted in Fig. 3, where its agreement with the QCD data is clear. For the strong coupling $g_{1}$, we find
$g_{1}\equiv\mathcal{G}_{1}(-m_{J/\psi}^{2})=(2.9\pm 0.5)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (26)
Partial width of the decay $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$ is given by the formula
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon\right]=g_{1}^{2}\frac{\lambda_{1}}{40\pi m^{2}}|M_{1}|^{2},$ (27)
where
$|M_{1}|^{2}=\frac{1}{6m^{4}}\left[m_{J/\psi}^{8}+m_{J/\psi}^{6}(m^{2}-4m_{\Upsilon}^{2})+(m^{2}-m_{\Upsilon}^{2})^{2}\right$.
$\times(6m^{4}+3m^{2}m_{\Upsilon}^{2}+m_{\Upsilon}^{4})+m_{J/\psi}^{4}(m^{4}-m^{2}m_{\Upsilon}^{2}+6m_{\Upsilon}^{4})$
$\left.-m_{J/\psi}^{2}(9m^{6}-34m^{4}m_{\Upsilon}^{2}+m^{2}m_{\Upsilon}^{4}+4m_{\Upsilon}^{6})\right].$ (28)
In Eq. (27), $\lambda_{1}=\lambda(m,m_{\Upsilon},m_{J/\psi})$ is determined by means of the expression
$\lambda(x,y,z)=\frac{\sqrt{x^{4}+y^{4}+z^{4}-2(x^{2}y^{2}+x^{2}z^{2}+y^{2}z^{2})}}{2x}.$ (29)
Finally, one gets
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon\right]=(23.7\pm 6.8)~\mathrm{MeV}.$ (30)
The errors originate from the ambiguities in the strong coupling $g_{1}$ and those of the particles’ masses in Eq. (27)
Figure 3: QCD data and fit function for $g_{1}(Q^{2})$. The diamond fixes the point $Q^{2}=-m_{J/\psi}^{2}$ where $g_{1}$ has been estimated.
III.2 $\mathcal{M}_{\mathrm{T}}\rightarrow\eta_{b}\eta_{c}$
Parameters of this decay depend on the strong coupling $g_{2}$ at the vertex $\mathcal{M}_{\mathrm{T}}\eta_{b}\eta_{c}$. To find the relevant form factor $g_{2}(q^{2})$, one has to consider the correlator
$\Pi_{\mu\nu}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{\ J^{\eta_{b}}(y)$ (31)
$\times J^{\eta_{c}}(0)J_{\mu\nu}^{\dagger}(x)\}|0\rangle.$
The interpolating currents of the quarkonia $\eta_{c}$ and $\eta_{b}$ are
$J^{\eta_{c}}(x)=\overline{c}_{i}(x)i\gamma_{5}c_{i}(x),\ J^{\eta_{b}}(x)=\overline{b}_{j}(x)i\gamma_{5}b_{j}(x).$ (32)
To calculate $\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ it is necessary to introduce the matrix elements
$\langle 0|J^{\eta_{b}}|\eta_{b}(p^{\prime})\rangle=\frac{f_{\eta_{b}}m_{\eta_{b}}^{2}}{2m_{b}},$
$\langle 0|J^{\eta_{c}}|\eta_{c}(q)\rangle=\frac{f_{\eta_{c}}m_{\eta_{c}}^{2}}{2m_{c}}.$ (33)
Here, $f_{\eta_{b}}$, $m_{\eta_{b}}$ and $f_{\eta_{c}}$, $m_{\eta_{c}}$ are the decay constants and masses of the mesons $\eta_{b}$ and $\eta_{c}$, respectively. The vertex $\mathcal{M}_{\mathrm{T}}\eta_{b}\eta_{c}$ is determined by the expression Agaev:2024pil
$\langle\eta_{b}(p^{\prime})\eta_{c}(q)|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle=g_{2}(q^{2})\epsilon_{\alpha\beta}^{(\lambda)}(p)p^{\prime\alpha}p^{\prime\beta}.$ (34)
Then $\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ is
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=g_{2}(q^{2})\frac{\Lambda f_{\eta_{b}}m_{\eta_{b}}^{2}f_{\eta_{c}}m_{\eta_{c}}^{2}}{4m_{b}m_{c}\left(p^{2}-m^{2}\right)(p^{\prime 2}-m_{\eta_{b}}^{2})}$
$\times\frac{1}{(q^{2}-m_{\eta_{c}}^{2})}\left[\frac{m^{4}-2m^{2}(m_{\eta_{b}}^{2}+q^{2})+(m_{\eta_{b}}^{2}-q^{2})^{2}}{12m^{2}}g_{\mu\nu}\right$.
$\left.+p_{\mu}^{\prime}p_{\nu}^{\prime}+\text{other terms}\right].$ (35)
The function $\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$ reads
$\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=i\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{5}S_{b}^{ib}(y-x)\right$.
$\left.\times\gamma_{\mu}S_{c}^{bj}(x)\gamma_{5}S_{c}^{ja}(-x)\gamma_{\nu}S_{b}^{ai}(x-y)\right].$ (36)
The terms in $\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ and $\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$ have the identical Lorentz compositions. To derive the sum rule for $g_{2}(q^{2})$, we choose components $\sim p_{\mu}^{\prime}p_{\nu}^{\prime}$ and utilize related amplitudes $\Pi_{2}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ and $\Pi_{2}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$. As a result, we get
$g_{2}(q^{2})=\frac{4m_{c}m_{b}(q^{2}-m_{\eta_{c}}^{2})}{\Lambda f_{\eta_{c}}m_{\eta_{c}}^{2}f_{\eta_{b}}m_{\eta_{b}}^{2}}e^{m^{2}/M_{1}^{2}}e^{m_{\eta_{b}}^{2}/M_{2}^{2}}$
$\times\Pi_{2}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}),$ (37)
where $\Pi_{2}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ is the amplitude $\Pi_{2}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ after relevant transformations and subtractions.
Operations to extract $g_{2}(q^{2})$ have been detailed above and are standard recipes of the SR method. In numerical analysis, we use $m_{\eta_{c}}=(2984.1\pm 0.4)~\mathrm{MeV}$, $m_{\eta_{b}}=(9398.7~\pm 2.0)\ \mathrm{MeV}$ as the masses of the mesons $\eta_{c}$ and $\eta_{b}$ PDG:2024 . We employ also the decay constants $f_{\eta_{c}}=(421\pm 35)~\mathrm{MeV}$, and $f_{\eta_{b}}=724~\mathrm{MeV}$, where the former is the sum rule prediction Veliev:2010vd . For the parameters $M_{2}^{2}$, and $s_{0}^{\prime}$ in the $\eta_{b}$ channel, we have applied the working windows
$M_{2}^{2}\in[10,12]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[95,99]~\mathrm{GeV}^{2}.$ (38)
The extrapolating function $\mathcal{G}_{2}(Q^{2})$ with parameters $\mathcal{G}_{2}^{0}=16.53~\mathrm{GeV}^{-1}$, $c_{2}^{1}=8.45$, and $c_{2}^{2}=-14.10$ nicely agrees with QCD predictions. The strong coupling $g_{2}$ amounts to
$g_{2}\equiv\mathcal{G}_{2}(-m_{\eta_{c}}^{2})=(10.07\pm 1.91)\ \mathrm{GeV}^{-1}.$ (39)
The partial width of this process is equal to
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow\eta_{b}\eta_{c}\right]=g_{2}^{2}\frac{\lambda_{2}}{40\pi m^{2}}|M_{2}|^{2},$ (40)
where
$|M_{2}|^{2}=\frac{\left[m^{4}+(m_{\eta_{b}}^{2}-m_{\eta_{c}}^{2})^{2}-2m^{2}(m_{\eta_{b}}^{2}+m_{\eta_{c}}^{2})\right]^{2}}{24m^{4}},$
and $\lambda_{2}=\lambda(m,m_{\eta_{b}},m_{\eta_{c}})$.
The decay $\mathcal{M}_{\mathrm{T}}\rightarrow\eta_{b}\eta_{c}$ has the width
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow\eta_{b}\eta_{c}\right]=(18.3\pm 8.6)~\mathrm{MeV}.$ (42)
IV Decays $\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{\ast+}B_{c}^{\ast-}$ and $B_{c}^{+}B_{c}^{-}$
In this section, we explore decays of the tensor molecule $\mathcal{M}_{\mathrm{T}}$ to $B_{c}^{\ast+}B_{c}^{\ast-}$ and $B_{c}^{+}B_{c}^{-}$ mesons. It is easy to see that the decays $\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{\ast+}B_{c}^{\ast-}$ and $B_{c}^{+}B_{c}^{-}$ with thresholds $12.68~\mathrm{GeV}$ and $12.55~\mathrm{GeV}$ are kinematically allowed modes of the tensor molecule $\mathcal{M}_{\mathrm{T}}$.
For the masses of these mesons, we are going to employ
$m_{B_{c}}$ $=$ $(6274.47\pm 0.27\pm 0.17)~\mathrm{MeV},$
$m_{B_{c}^{\ast}}$ $=$ $6338~\mathrm{MeV},$ (43)
where first of them is the experimental datum PDG:2024 , whereas the second one is the theoretical prediction from Ref. Godfrey:2004ya . The decay constant of these mesons are
$f_{B_{c}}=(371\pm 37)~\mathrm{MeV},\ f_{B_{c}^{\ast}}=471~\mathrm{MeV},$ (44)
which are borrowed from Refs. Wang:2024fwc ; Eichten:2019gig , respectively.
IV.1 $\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{\ast+}B_{c}^{\ast-}$
Analysis of this decay is done in accordance with the scheme presented above. Therefore, we provide only decisive formulas and predictions.
Our aim is to extract SR for the form factor $g_{3}(q^{2})$ which describe the interaction of particles at the vertex $\mathcal{M}_{\mathrm{T}}B_{c}^{\ast+}B_{c}^{\ast-}$. The relevant correlator is
$\widetilde{\Pi}_{\mu\nu\alpha\beta}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{B_{c}^{\ast+}}(y)$ (45)
$\times J_{\nu}^{B_{c}^{\ast-}}(0)J_{\alpha\beta}^{\dagger}(x)\}|0\rangle.$
Here, $J_{\mu}^{B_{c}^{\ast+}}$ and $J_{\nu}^{B_{c}^{\ast-}}$ are the currents that correspond to $B_{c}^{\ast+}$ and $B_{c}^{\ast-}$ mesons
$J_{\mu}^{B_{c}^{\ast+}}(x)=\overline{b}_{i}(x)\gamma_{\mu}c_{i}(x),\ J_{\nu}^{B_{c}^{\ast-}}(x)=\overline{c}_{j}(x)\gamma_{\nu}b_{j}(x).$ (46)
After some operations the correlation function $\widetilde{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$ acquires the form
$\widetilde{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{B_{c}^{\ast+}}|B_{c}^{\ast+}(p^{\prime},\varepsilon_{1})\rangle}{p^{\prime 2}-m_{B_{c}^{\ast}}^{2}}\frac{\langle 0|J_{\nu}^{B_{c}^{\ast-}}|B_{c}^{\ast-}(q,\varepsilon_{2})\rangle}{q^{2}-m_{B_{c}^{\ast}}^{2}}$
$\times\langle B_{c}^{\ast+}(p^{\prime},\varepsilon_{1})B_{c}^{\ast-}(q,\varepsilon_{2})|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle\frac{\langle\mathcal{M}_{\mathrm{T}}(p,\epsilon)|J_{\alpha\beta}^{\dagger}|0\rangle}{p^{2}-m^{2}}$
$+\cdots.$ (47)
Subsequent calculations are performed using the matrix elements
$\langle 0|J_{\mu}^{B_{c}^{\ast+}}|B_{c}^{\ast+}(p^{\prime},\varepsilon_{1})\rangle=f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}\varepsilon_{1\mu}(p^{\prime}),$
$\langle 0|J_{\nu}^{B_{c}^{\ast-}}|B_{c}^{\ast-}(q,\varepsilon_{2})\rangle=f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}\varepsilon_{2\nu}(q),$ (48)
where $\varepsilon_{1\mu}(p^{\prime})$ and $\varepsilon_{2\nu}(q)$ are the polarization vectors of $B_{c}^{\ast+}$ and $B_{c}^{\ast-}$, respectively. The vertex $\mathcal{M}_{\mathrm{T}}B_{c}^{\ast+}B_{c}^{\ast-}$ is treated as in Eq. (18) with substitution $g_{1}(q^{2})\rightarrow g_{3}(q^{2})$.
Then, the correlation function $\widetilde{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$ in terms of the parameters of the molecule $\mathcal{M}_{\mathrm{T}}$ and $B_{c}^{\ast\pm}$ mesons is
$\widetilde{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})=\frac{g_{3}(q^{2})\Lambda f_{B_{c}^{\ast}}^{2}m_{B_{c}^{\ast}}^{2}}{\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{B_{c}^{\ast}}^{2}\right)}$
$\times\frac{1}{(q^{2}-m_{B_{c}^{\ast}}^{2})}\left[\frac{1}{2}p_{\mu}p_{\alpha}^{\prime}g_{\nu\beta}+\frac{m^{2}+m_{B_{c}^{\ast}}^{2}-q^{2}}{4m^{2}}p_{\mu}^{\prime}p_{\alpha}g_{\beta\nu}\right$.
$\left.+p_{\beta}^{\prime}p_{\alpha}^{\prime}g_{\mu\nu}+\text{other structures}\right]+\cdots.$ (49)
The function $\widetilde{\Pi}_{\mu\nu\alpha\beta}(p,p^{\prime})$ computed using quark propagators amounts to
$\widetilde{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p,p^{\prime})=i\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{ia}(y-x)\right$.
$\left.\times\gamma_{\beta}S_{b}^{ai}(x-y)\right]\mathrm{Tr}\left[\gamma_{\nu}S_{b}^{jb}(-x)\gamma_{\alpha}S_{c}^{bj}(x)\right].$
The sum rule for $g_{3}(q^{2})$
$g_{3}(q^{2})=\frac{2(q^{2}-m_{B_{c}^{\ast}}^{2})}{\Lambda f_{B_{c}^{\ast}}^{2}m_{B_{c}^{\ast}}^{2}}e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}^{\ast}}^{2}/M_{2}^{2}}$
$\times\Pi_{3}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ (51)
is extracted by employing the amplitudes $\Pi_{3}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ and $\Pi_{3}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ which correspond to terms $\sim p_{\mu}p_{\alpha}^{\prime}g_{\nu\beta}$ in the correlators $\widetilde{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$ and $\widetilde{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p,p^{\prime})$, respectively. Above $\Pi_{3}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ is the amplitude $\Pi_{3}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ obtained after relevant transformations.
[…]
$M_{2}^{2}\in[6.5,7.5]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[49,51]~\mathrm{GeV}^{2}.$ (52)
The constants of the function $\mathcal{G}_{3}(Q^{2})$ are equal to $\mathcal{G}_{3}^{0}=0.59~\mathrm{GeV}^{-1}$, $c_{3}^{1}=3.53$, and $c_{3}^{2}=-2.84$. For the coupling $g_{3}$, we obtain
$g_{3}\equiv\mathcal{G}_{3}(-m_{B_{c}^{\ast}}^{2})=(2.1\pm 0.4)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (53)
The partial width of the decay $\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{\ast+}B_{c}^{\ast-}$ is equal to
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{\ast+}B_{c}^{\ast-}\right]=\frac{g_{3}^{2}\lambda_{3}}{80\pi m^{2}}\left(m^{4}-3m^{2}m_{B_{c}^{\ast}}^{2}\right$.
$\left.+6m_{B_{c}^{\ast}}^{4}\right),$ (54)
where $\lambda_{3}=\lambda(m,m_{B_{c}^{\ast}},m_{B_{c}^{\ast}})$.
Numerical calculations yield
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{\ast+}B_{c}^{\ast-}\right]=(20.7\pm 6.9)~\mathrm{MeV}.$ (55)
IV.2 $\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{+}B_{c}^{-}$
The decay $\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{+}B_{c}^{-}$ is studied in the same way. We consider the correlation function
$\widetilde{\Pi}_{\mu\nu}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{\ J^{B_{c}^{+}}(y)$ (56)
$\times J^{B_{c}^{-}}(0)J_{\mu\nu}^{\dagger}(x)\}|0\rangle,$
with $J^{B_{c}^{+}}(x)$ and $J^{B_{c}^{-}}(x)$ being the interpolating currents of the $B_{c}^{+}$ and $B_{c}^{-}$ mesons
$J^{B_{c}^{+}}(x)=\overline{b}_{i}(x)i\gamma_{5}c_{i}(x),\ J^{B_{c}^{-}}(x)=\overline{c}_{j}(x)i\gamma_{5}b_{j}(x).$ (57)
The matrix elements of the pseudoscalar $B_{c}^{\pm}$ mesons are
$\langle 0|J^{B_{c}^{\pm}}|B_{c}^{\pm}\rangle=\frac{f_{B_{c}}m_{B_{c}}^{2}}{m_{b}+m_{c}}.$ (58)
The vertex $\langle B_{c}^{+}(p^{\prime})B_{c}^{-}(q)|T(p,\epsilon)\rangle$ has the form
$\langle B_{c}^{+}(p^{\prime})B_{c}^{-}(q)|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle=g_{4}(q^{2})\epsilon_{\alpha\beta}^{(\lambda)}(p)p^{\prime\alpha}p^{\prime\beta}.$ (59)
The $\widetilde{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ is given after some substitutions by Eq. (35), while Eq. (36) determines the QCD side of SR. The form factor $g_{4}(q^{2})$ is
$g_{4}(q^{2})=\frac{(m_{b}+m_{c})^{2}(q^{2}-m_{B_{c}}^{2})}{\Lambda f_{B_{c}}^{2}m_{B_{c}}^{4}}e^{m^{2}/M_{1}^{2}}e^{m_{B_{c}}^{2}/M_{2}^{2}}$
$\times\Pi_{4}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (60)
In calculations, we have utilized the parameters
$M_{2}^{2}\in[6.5,7.5]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[45,47]~\mathrm{GeV}^{2}.$ (61)
Computations of the form factor $g_{4}(q^{2})$ and coupling $g_{4}$ lead to the result
$g_{4}\equiv\mathcal{G}_{4}(-m_{B_{c}}^{2})=(10.67\pm 1.92)\ \mathrm{GeV}^{-1},$ (62)
where $\mathcal{G}_{4}(Q^{2})$ is the fitting function with parameters $\mathcal{G}_{4}^{0}=14.62~\mathrm{GeV}^{-1}$, $c_{4}^{1}=3.09$, and $c_{4}^{2}=7.42$.
The width of the decay $\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{+}B_{c}^{-}$ is calculated by utilizing the formula
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{+}B_{c}^{-}\right]$ $=$ $g_{4}^{2}\frac{\lambda_{4}}{960\pi m^{2}}\left(m^{2}-4m_{B_{c}}^{2}\right)^{2},$
where $\lambda_{4}=\lambda(m,m_{B_{c}},m_{B_{c}})$. This formula can be obtained from Eq. (41) in the limit $m_{\eta_{b}}$ and $m_{\eta_{c}}\rightarrow m_{B_{c}}$ . Our computations yield
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow B_{c}^{+}B_{c}^{-}\right]=(21.2\pm 12.3)~\mathrm{MeV}.$ (64)
V Decays generated by $b\overline{b}$ annihilations
The hadronic molecule $\mathcal{M}_{\mathrm{T}}$ can be converted to ordinary mesons also after annihilation of $b\overline{b}$ quarks to light quark-antiquark pairs Becchi:2020mjz ; Becchi:2020uvq ; Agaev:2023ara and generation of $DD$ mesons of appropriate parameters. We study here the decays of the molecule $\mathcal{M}_{\mathrm{T}}$ to $D^{+}D^{-}$, $^{0}\overline{D}^{0}, D+D$$^{\ast+}D^{\ast-}$, $D^{\ast 0}\overline{D}^{\ast 0}$, $D_{s}^{+}D_{s}^{-}$ and $D_{s}^{\ast+}D_{s}^{\ast-}$ mesons.
These modes are kinematically possible channels of the molecule $\mathcal{M}_{\mathrm{T}}$ to conventional particles. We investigate these decays also in the three-point SR framework. Here, to calculate the correlation functions which contain vacuum matrix element $\langle\overline{b}b\rangle$ of $b$ quark, we replace it with the known gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ Agaev:2023ara .
V.1 Decays $\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast 0}\overline{D}^{\ast 0}$ and $D^{\ast+}D^{\ast-}$
First, we analyze the process $\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast 0}\overline{D}^{\ast 0}$. To find the coupling $G_{1}$ of particles at the vertex $\mathcal{M}_{\mathrm{T}}D^{\ast 0}\overline{D}^{\ast 0}$, we consider the correlator
$\widehat{\Pi}_{\mu\nu\alpha\beta}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{\overline{D}^{\ast 0}}(y)$ (65)
$\times J_{\nu}^{D^{\ast 0}}(0)J_{\alpha\beta}^{\dagger}(x)\}|0\rangle,$
with $J_{\mu}^{\overline{D}^{\ast 0}}(x)$ and $J_{\nu}^{D^{\ast 0}}(x)$ being the interpolating currents for the mesons $\overline{D}^{\ast 0}$ and $D^{\ast 0}$
$J_{\mu}^{\overline{D}^{\ast 0}}(x)=\overline{c}_{i}(x)\gamma_{\mu}u_{i}(x),\text{ }J_{\nu}^{D^{\ast 0}}(x)=\overline{u}_{j}(x)\gamma_{\nu}c_{j}(x).$ (66)
Figure 4: QCD data and extrapolating functions for the form factors $G_{1}(Q^{2})$ (dashed curve) and $G_{3}(Q^{2})$ (solid curve). The circle and diamond fix the points $Q^{2}=-m_{D^{\ast 0}}^{2}$ and $Q^{2}=-m_{D^{0}}^{2}$, respectively.
The correlation function $\widehat{\Pi}_{\mu\nu\alpha\beta}(p,p^{\prime})$ in terms of the matrix elements of the particles $\mathcal{M}_{\mathrm{T}}$, $D^{\ast 0}$, and $\overline{D}^{\ast 0}$ is
$\widehat{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{\overline{D}^{\ast 0}}|\overline{D}^{\ast 0}(p^{\prime},\varepsilon_{1})\rangle}{p^{\prime 2}-m_{D^{\ast 0}}^{2}}\frac{\langle 0|J_{\nu}^{D^{\ast 0}}|D^{\ast 0}(q,\varepsilon_{2})\rangle}{q^{2}-m_{D^{\ast 0}}^{2}}$
$\times\langle\overline{D}^{\ast 0}(p^{\prime},\varepsilon_{1})D^{\ast 0}(q,\varepsilon_{2})|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle\frac{\langle\mathcal{M}_{\mathrm{T}}(p,\epsilon)|J_{\mu}^{\dagger}|0\rangle}{p^{2}-m^{2}}$
$+\cdots,$ (67)
where $m_{D^{\ast 0}}=(2006.85\pm 0.05)~\mathrm{MeV}$ is the mass of the mesons $\overline{D}^{\ast 0}$ and $D^{\ast 0}$, whereas $\varepsilon_{1\mu}$ and $\varepsilon_{2\nu}$ are their polarization vectors.
The function $\widehat{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$ is found by means of the matrix elements
$\langle 0|J_{\mu}^{\overline{D}^{\ast 0}}|\overline{D}^{\ast 0}(p^{\prime},\varepsilon_{1})\rangle$ $=$ $f_{D^{\ast}}m_{D^{\ast 0}}\varepsilon_{1\mu}(p^{\prime}),$
$\langle 0|J_{\nu}^{D^{\ast 0}}|D^{\ast 0}(q,\varepsilon_{2})\rangle$ $=$ $f_{D^{\ast}}m_{D^{\ast 0}}\varepsilon_{2\nu}(q),$ (68)
with $f_{D^{\ast}}=(252.2\pm 22.66)~\mathrm{MeV}$ being the decay constant of the mesons $D^{\ast 0}$ and $\overline{D}_{D}^{\ast 0}$. The vertex $\langle\overline{D}^{\ast 0}(p^{\prime},\varepsilon_{1})D^{\ast 0}(q,\varepsilon_{2})|\mathcal{M}_{\mathrm{T}}(p,\epsilon)\rangle$ is written down in the form of Eq. (18).
The correlation function $\widehat{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$ contains different components. The sum rule for the form factor $G_{1}(q^{2})$ is obtained by employing the amplitude $\widehat{\Pi}_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ corresponding to the term $\sim p_{\mu}p_{\alpha}^{\prime}g_{\beta\nu}$. The correlator $\widehat{\Pi}_{\mu\nu\alpha\beta}(p,p^{\prime})$ calculated by employing the quark propagators amounts to
$\widehat{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\langle\overline{b}b\rangle\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}$
$\times\mathrm{Tr}\left[\gamma_{\mu}S_{c}^{ja}(y-x)\gamma_{\beta}\gamma_{\alpha}S_{c}^{ai}(x)\gamma_{\nu}{}S_{u}^{ij}(-y)\right],$ (69)
where $S_{u}(x)$ is the $u$ quark’s propagator Agaev:2020zad . We denote by $\widehat{\Pi}_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ the amplitude that in $\widehat{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p,p^{\prime})$ corresponds to the same term $\sim p_{\mu}p_{\alpha}^{\prime}g_{\beta\nu}$.
To continue computations, we utilize the relation between the condensates
$\langle\overline{b}b\rangle=-\frac{1}{12m_{b}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle$ (70)
obtained in Ref. Shifman:1978bx by employing the sum rules.
The sum rule for the form factor $G_{1}(q^{2})$ is
$G_{1}(q^{2})=\frac{2(q^{2}-m_{D^{\ast 0}}^{2})}{\Lambda f_{D^{\ast}}^{2}m_{D^{\ast 0}}^{2}}e^{m^{2}/M_{1}^{2}}e^{m_{D^{\ast 0}}^{2}/M_{2}^{2}}$
$\times\widehat{\Pi}_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}),$ (71)
where $\widehat{\Pi}_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ is the amplitude $\widehat{\Pi}_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ after the Borel transformations and continuum subtractions.
In numerical computations for the $\overline{D}^{\ast 0}$ meson channel, we utilize the parameters
$M_{2}^{2}\in[2,3]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[5.7,5.8]~\mathrm{GeV}^{2}.$ (72)
To evaluate the coupling $G_{1}$ we use the SR data for $Q^{2}=2-30\ \mathrm{GeV}^{2}$ and the extrapolating function with parameters $\widehat{\mathcal{G}}_{1}^{0}=0.056~\mathrm{GeV}^{-1}$, $\widehat{c}_{1}^{1}=8.32$, and $\widehat{c}_{1}^{2}=-14.87$. The sum rule data and function $\widehat{\mathcal{G}}_{1}(Q^{2})$ are plotted in Fig. 4. The coupling $G_{1}$ is calculated at $q^{2}=m_{D^{\ast 0}}^{2}$
$G_{1}\equiv\widehat{\mathcal{G}}_{1}(-m_{D^{\ast 0}}^{2})=(4.5\pm 0.9)\times 10^{-2}\ \mathrm{GeV}^{-1}.$ (73)
The width of the decay $\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast 0}\overline{D}^{\ast 0}$ is
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast 0}\overline{D}^{\ast 0}\right]=(7.63\pm 2.17)~\mathrm{MeV}.$ (74)
The difference between the second $\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast+}D^{\ast-}$ and first decays is encoded in the masses of the final-state mesons. The QCD side of the channel $\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast+}D^{\ast-}$ can be obtained from Eq. (69) by the substitution $S_{u}^{ij}(-y)\rightarrow S_{d}^{ij}(-y)$. Because in this paper we adopt the equality $m_{u}=m_{d}=0$ this means that QCD sides of these two decays are identical to each other. Therefore, with nice accuracy $\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast+}D^{\ast-}\right]\approx\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast 0}\overline{D}^{\ast 0}\right]$. Nevertheless, we have carried out corresponding computations and found
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D^{\ast+}D^{\ast-}\right]=(7.67\pm 2.17)~\mathrm{MeV}.$ (75)
As it has been expected, the difference between partial widths of these channels is very small.
V.2 Processes $\mathcal{M}_{\mathrm{T}}\rightarrow D^{0}\overline{D}^{0}$ and $D^{+}D^{-}$
The channel $\mathcal{M}_{\mathrm{T}}\rightarrow D^{0}\overline{D}^{0}$ is considered using the correlation function
$\widehat{\Pi}_{\mu\nu}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J^{\overline{D}^{0}}(y)$ (76)
$\times J^{D^{0}}(0)J_{\mu\nu}^{\dagger}(x)\}|0\rangle,$
where the currents $J^{\overline{D}^{0}}(x)$ and $J^{D^{0}}(x)$ are introduced by formulas
$J^{\overline{D}^{0}}(x)=\overline{c}_{i}(x)i\gamma_{5}u_{i}(x),\text{ }J^{D^{0}}(x)=\overline{u}_{j}(x)i\gamma_{5}c_{j}(x).$ (77)
To derive the sum rule for the form factor $G_{3}(q^{2})$ that describes the strong interaction of particles at the vertex $\mathcal{M}_{\mathrm{T}}D^{0}\overline{D}^{0}$, we have to find the correlators $\widehat{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ and $\widehat{\Pi}_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$.
We determine $\widehat{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ by means of the matrix elements
$\langle 0|J^{\overline{D}^{0}}|\overline{D}^{0}\rangle=\langle 0|J^{D^{0}}|D^{0}\rangle=\frac{f_{D}m_{D^{0}}^{2}}{m_{c}},$ (78)
and
$\langle\overline{D}^{0}(p^{\prime})D^{0}(q)|T(p,\epsilon)\rangle=G_{3}(q^{2})\epsilon_{\alpha\beta}^{(\lambda)}(p)p^{\prime\alpha}p^{\prime\beta},$ (79)
where $m_{D^{0}}=(1864.84\pm 0.05)~\mathrm{MeV}$ and $f_{D}=(211.9\pm 1.1)~\mathrm{MeV}$ are the mass and decay constant of mesons $D^{0}$ and $\overline{D}^{0}$ PDG:2024 ; Rosner:2015wva . As a result, we find
$\widehat{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})=\frac{G_{3}(q^{2})\Lambda f_{D}^{2}m_{D^{0}}^{4}}{m_{c}^{2}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{D^{0}}^{2}\right)\left(q^{2}-m_{D^{0}}^{2}\right)}$
$\times\left[\frac{m^{4}-2m^{2}(m_{D^{0}}^{2}+q^{2})+(m_{D^{0}}^{2}-q^{2})^{2}}{12m^{2}}g_{\mu\nu}\right$.
$\left.+p_{\mu}^{\prime}p_{\nu}^{\prime}-\frac{m^{2}+m_{D^{0}}^{2}-q^{2}}{2m^{2}}p_{\mu}p_{\nu}^{\prime}+\text{other terms}\right].$ (80)
For $\widehat{\Pi}_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$, we get
$\widehat{\Pi}_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\langle\overline{b}b\rangle\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}$
$\times\mathrm{Tr}\left[\gamma_{5}{}S_{c}^{ja}(y-x)\gamma_{\mu}\gamma_{\nu}S_{c}^{ai}(x)\gamma_{5}S_{u}^{ij}(-y)\right].$ (81)
To obtain the sum rule for $G_{3}(q^{2})$, we employ the amplitudes $\widehat{\Pi}_{3}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ and $\widehat{\Pi}_{3}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ corresponding to the structure $p_{\mu}^{\prime}p_{\nu}^{\prime}$ and find
$G_{3}(q^{2})=\frac{m_{c}^{2}(q^{2}-m_{D^{0}}^{2})}{\Lambda f_{D}^{2}m_{D^{0}}^{4}}e^{m^{2}/M_{1}^{2}}e^{m_{D^{0}}^{2}/M_{2}^{2}}$
$\times\widehat{\Pi}_{3}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}),$ (82)
with $\widehat{\Pi}_{3}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ being the transformed function $\widehat{\Pi}_{3}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$.
In numerical calculations we have used the parameters
$M_{2}^{2}\in[1.5,3]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[5,5.2]~\mathrm{GeV}^{2}.$ (83)
The coupling $G_{3}$ is estimated by means of the function $\widehat{\mathcal{G}}_{3}(Q^{2})$ with $\widehat{\mathcal{G}}_{3}^{0}=0.20~\mathrm{GeV}^{-1}$, $\widehat{c}_{3}^{1}=11.02$, and $\widehat{c}_{3}^{2}=-28.26$
$G_{3}\equiv\widehat{\mathcal{G}}_{3}(-m_{D^{0}}^{2})=(1.53\pm 0.28)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (84)
The partial width of the decay $\mathcal{M}_{\mathrm{T}}\rightarrow D^{0}\overline{D}^{0}$ is equal to
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D^{0}\overline{D}^{0}\right]=(6.64\pm 1.73)~\mathrm{MeV}.$ (85)
The width of the process $\mathcal{M}_{\mathrm{T}}\rightarrow D^{+}D^{-}$, as it has been explained above, is approximately equal to $\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D^{0}\overline{D}^{0}\right]$. Indeed, our computations yield
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D^{+}D^{-}\right]=(6.45\pm 1.64)~\mathrm{MeV}.$ (86)
V.3 Channels $\mathcal{M}_{\mathrm{T}}\rightarrow D_{s}^{+}D_{s}^{-}$ and $D_{s}^{\ast+}D_{s}^{\ast-}$
Investigation of these channels does not differ considerably from analysis of the modes considered in the previous subsections. There are only a few replacements in the correlation functions and parameters of the mesons which should be taken into account.
Thus, the correlators of the decays $\mathcal{M}_{\mathrm{T}}\rightarrow D_{s}^{\ast+}D_{s}^{\ast-}$ and $\mathcal{M}_{\mathrm{T}}\rightarrow D_{s}^{+}D_{s}^{-}$ are obtained from Eqs. (69) and (81) after replacing the propagators $S_{u}^{ij}(-y)$ by $S_{s}^{ij}(-y)$. In numerical computations we include in analysis terms $\sim m_{s}=(93.5\pm 0.8)~\mathrm{MeV}$, but ignore ones proportional to $m_{s}^{2}$. It is clear that the masses and decay constants of the $D_{s}$ and $D_{s}^{\ast}$ mesons
$m_{D_{s}}$ $=$ $(1969.0\pm 1.4)~\mathrm{MeV},\ f_{D_{s}}=(249.9\pm 0.5)~\mathrm{MeV},$
and
$m_{D_{s}^{\ast}}$ $=$ $(2112.2\pm 0.4)~\mathrm{MeV},\ f_{D_{s}^{\ast}}=(268.8\pm 6.5)~\mathrm{MeV},$
appear in calculations as input parameters.
The widths of these decays are
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D_{s}^{+}D_{s}^{-}\right]=(6.34\pm 1.72)~\mathrm{MeV},$ (89)
and
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow D_{s}^{\ast+}D_{s}^{\ast-}\right]=(7.38\pm 2.10)~\mathrm{MeV},$ (90)
respectively.
VI Decays into $B$ meson pairs
Decays of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$ into pairs of the $B$ and $B_{s}$ mesons are investigated by the same manner. These decays are generated by annihilation in the molecule of $c\overline{c}$ quarks which lead to creations of $BB$ and $B_{s}B_{s}$ mesons with correct charges and quantum numbers. It is easy to see that $\mathcal{M}_{\mathrm{T}}\rightarrow B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{0}\overline{B}_{s}^{0}$ are allowed decay channels of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$.
Below, as a sample, we consider the decay $\mathcal{M}_{\mathrm{T}}\rightarrow B^{\ast+}B^{\ast-}$. The correlation function necessary to evaluate the coupling $\overline{G}_{1}$ at the vertex $\mathcal{M}_{\mathrm{T}}B^{\ast+}B^{\ast-}$ is
$\overline{\Pi}_{\mu\nu\alpha\beta}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{B^{\ast+}}(y)$ (91)
$\times J_{\nu}^{B^{\ast-}}(0)J_{\alpha\beta}^{\dagger}(x)\}|0\rangle,$
where $J^{B^{\ast+}}(x)$ and $J^{B^{\ast-}}(x)$ are the interpolating currents of the mesons $B^{\ast+}$ and $B^{\ast-}$
$J_{\mu}^{B^{\ast+}}(x)=\overline{b}_{i}(x)\gamma_{\mu}u_{i}(x),\ J_{\nu}^{B^{\ast-}}(x)=\overline{u}_{j}(x)\gamma_{\nu}b_{j}(x).$ (92)
In calculation of the correlator $\overline{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{Phys}}(p,p^{\prime})$, we have used the matrix elements
$\langle 0|J_{\mu}^{B^{\ast+}}|B^{\ast+}(p^{\prime},\varepsilon_{1})\rangle=f_{B^{\ast}}m_{B^{\ast}}\varepsilon_{1\mu}(p^{\prime}),$
$\langle 0|J_{\nu}^{B^{\ast-}}|B^{\ast-}(q,\varepsilon_{2})\rangle=f_{B^{\ast}}m_{B^{\ast}}\varepsilon_{2\nu}(q).$ (93)
In these expressions $m_{B^{\ast}}=(5324.75\pm 0.20)~\mathrm{MeV}$ and $f_{B^{\ast}}=(210\pm 6)~\mathrm{MeV}$ are the mass and decay constant of the $B^{\ast\pm}$ mesons. The matrix element of the vertex $\mathcal{M}_{\mathrm{T}}B^{\ast+}B^{\ast-}$ is given by Eq. (18) after evident replacements.
The correlator $\overline{\Pi}_{\mu\nu\alpha\beta}(p,p^{\prime})$ through quark propagators can be expressed by the following formula
$\overline{\Pi}_{\mu\nu\alpha\beta}^{\mathrm{OPE}}(p,p^{\prime})=\frac{1}{3}\langle\overline{c}c\rangle\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}$
$\times\mathrm{Tr}\left[\gamma_{\mu}S_{u}^{ij}(y)\gamma_{\nu}S_{b}^{ja}(-x)\gamma_{\alpha}\gamma_{\beta}S_{b}^{ai}(x-y){}\right].$ (94)
The condensate $\langle\overline{c}c\rangle$ with some accuracy can be replaced by Shifman:1978bx
$\langle\overline{c}c\rangle=-\frac{1}{12m_{c}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle.$ (95)
Numerical calculations of the form factor $\overline{G}_{1}(q^{2})$ are performed for $Q^{2}=2-30~\mathrm{GeV}^{2}$. The Borel and continuum subtraction parameters in the $B^{\ast+}$ meson channel are varied within limits
$M_{2}^{2}\in[5.5,6.5]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[34,35]~\mathrm{GeV}^{2}.$ (96)
The coupling $\overline{G}_{1}$ is extracted at $Q^{2}=-m_{B^{\ast}}^{2}$ by utilizing the fit function $\overline{\mathcal{G}}_{1}(Q^{2})$ with parameters $\overline{\mathcal{G}}_{1}^{0}=0.104~\mathrm{GeV}^{-1}$, $\overline{c}_{1}^{1}=2.71$, and $\overline{c}_{1}^{2}=-3.63$
$\overline{G}_{1}\equiv\overline{\mathcal{G}}_{1}(-m_{B^{\ast}}^{2})=(5.89\pm 1.12)\times 10^{-2}\ \mathrm{GeV}^{-1}.$ (97)
Then the width of the process $\mathcal{M}_{\mathrm{T}}\rightarrow B^{\ast+}B^{\ast-}$ is equal to
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\rightarrow B^{\ast+}B^{\ast-}\right]=(5.46\pm 1.48)~\mathrm{MeV}.$ (98)
Another decays of $\mathcal{M}_{\mathrm{T}}$ to $B$ and $B_{s}$ mesons are studied in this manner. Results of relevant analyses are collected in Table 1, in which we give information on strong couplings and partial widths of these processes. Note that parameters of the decays $\mathcal{M}_{\mathrm{T}}\rightarrow B^{(\ast)0}\overline{B}^{(\ast)0}$ are chosen equal to those of $\mathcal{M}_{\mathrm{T}}\rightarrow B^{(\ast)+}B^{(\ast)-}$ and omitted in the Table.
Information on the partial widths of the channels considered in last three sections permit us to estimate the full width of the hadronic molecule
$\Gamma\left[\mathcal{M}_{\mathrm{T}}\right]=(154\pm 19)~\mathrm{MeV}.$ (99)
i Channels $\overline{G}_{i}~(\mathrm{GeV}^{-1})$ $\Gamma_{i}~(\mathrm{MeV})$
11 $B^{\ast+}B^{\ast-}$ $(5.89\pm 1.12)\times 10^{-2}$ $5.46\pm 1.48$
22 $B^{+}B^{-}$ $0.48\pm 0.09$ $4.98\pm 1.38$
33 $B_{s}^{\ast 0}\overline{B}_{s}^{\ast 0}$ $(5.07\pm 1.01)\times 10^{-2}$ $3.86\pm 1.10$
44 $B_{s}^{0}\overline{B}_{s}^{0}$ $0.44\pm 0.08$ $3.47\pm 0.94$
Table 1: Decay channels of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$ to $B$ and $B_{s}$ mesons, related couplings $\overline{G}_{i}$ and widths $\Gamma_{i}$.
VII Discussion and conclusions
In this article, we have computed important parameters of the hadronic tensor molecule $\mathcal{M}_{\mathrm{T}}=B_{c}^{\ast+}B_{c}^{\ast-}$ in the context of QCD sum rule method: We have evaluated its mass and full decay width. The mass of $\mathcal{M}_{\mathrm{T}}$ has been calculated by employing the two-point sum rule approach, whereas decays of these molecule have been investigated using technical tools of the three-point SR method.
In our previous publications Agaev:2025wdj ; Agaev:2025fwm , we explored parameters of the scalar and axial-vector molecules built of $B_{c}$ mesons. Comparing prediction for the masses of these states with $m=(12.87\pm 0.08)~\mathrm{GeV}$, we see that this particle is heavier than its counterparts. But this difference is relatively small and amounts to $\sim 100~\mathrm{MeV}$. It is interesting also to confront the tensor molecule $\mathcal{M}_{\mathrm{T}}$ with the tensor diquark-antidiquark state $T$ with the same quark content Agaev:2024qbh . The mass of the tetraquark $T$ was found there to be equal to $(12.70\pm 0.09)~\mathrm{GeV}$. As is seen, the gap between masses of these two structures is small and does not exceed $170~\mathrm{MeV}$, and even has a common point at $12.79~\mathrm{GeV}$ provided one takes into account uncertainties of computations.
The molecules $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$ were also considered in the framework of the coupled-channel unitary approach Liu:2023gla . In this article some integrals were regularized by including the parameter $\Lambda$. As a result, predictions for the masses of the molecules depend on a choice of $\Lambda$. For the mass of the tensor molecule $B_{c}^{\ast+}B_{c}^{\ast-}$ the maximal value $12660.7~\mathrm{MeV}$ was obtained at $\Lambda=400~\mathrm{MeV}$, whereas the minimal value $12287.7~\mathrm{MeV}$ was found at $\Lambda=1400~\mathrm{MeV}$. Theoretical ambiguities of these results are approximately $\sim 400~\mathrm{MeV}$. Even maximal value of the mass is smaller than our result. Note that the similar picture was observed while considering the scalar and axial-vector molecules as well.
Our prediction $m=(12.87\pm 0.08)~\mathrm{GeV}$ proves that it is unstable against the strong decays to different pairs of conventional mesons. The processes $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$, $\eta_{b}\eta_{c}$ and $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$ are main fall-apart channels of the molecule $\mathcal{M}_{\mathrm{T}}$: In these decays all heavy quarks from $\mathcal{M}_{\mathrm{T}}$ appear in the final-state mesons. Alternative modes of $\mathcal{M}_{\mathrm{T}}$ are due to annihilations of $b\overline{b}$ and $c\overline{c}$ quarks inside of the molecule and generation of various processes. Note that contribution of these decays to the full width of the molecule $\mathcal{M}_{\mathrm{T}}$ constitutes approximately a half of this parameter. Our result for $\Gamma\left[\mathcal{M}_{\mathrm{T}}\right]$ demonstrate that $\mathcal{M}_{\mathrm{T}}$ is a wide molecular structure, which can decay through two different mechanisms.
Four-quark mesons composed of heavy quarks are an import component of the exotic hadron spectroscopy. Hadronic molecules $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$ or diquark-antidiquark states with the same content $bc\overline{b}\overline{c}$ were not discovered, but they may be observed in the future LHC and FCC experiments Carvalho:2015nqf ; Abreu:2023wwg . Therefore, detailed theoretical studies of such systems by means of different approaches and comparison of the obtained predictions are necessary to prepare useful ground for these experiments.